Circular No. 5663

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SWIFT-TUTTLE (1992t) L. McFadden, University of Maryland, reports that spectroscopic ob-

servations (range 380-525 nm) of P/Swift-Tuttle made on Oct. 5.097 UT by D. Schleicher, R. Schultz, A. Chamberlin, and herself with the 1.8-m Perkins Telescope at Lowell Observatory (+ Ohio State CCD spectrograph) yield the following production rates, following a preliminary reduction by Schultz: log Q(CN) = 26.77, log Q(C_{3}) = 25.72, log Q(C_{2}) = 26.49.

M. A’Hearn, University of Maryland, communicates: “Observations of P/Swift-Tuttle by L. McFadden, P. Feldman, M. Haken, J. McPhate, and myself using the International Ultraviolet Explorer yield the following respective water production rates (standard vectorial model, in units of 10^{29} molecules/s): Nov. 3.1 UT, 1.5; 3.9, 2.2; 4.4, 2.5; 16.0, 3.4; 23.9 ∼ 5. CS production is roughly 0.2 percent, as in P/Halley, but higher than in most comets. Relative abundance can be significantly affected by phase lags during variability. The continuum brightness implies a gas-to-dust ratio comparable to but slightly higher than that of P/Halley. Photometric variability during the shifts indicates that periodic variations must have periods considerably greater than 16 hr.”

H. Mikuz, Ljubljana, Slovenia, reports that several 10-min CCD frames obtained around Nov. 19.78 and 20.75 UT, with a 250-mm f/4 lens and an H_{2}O^{+} 620-nm filter, show a straight ion tail ∼ 4^{o}.5 long in p.a. 41^{o}. Similar frames on Nov. 23.73-23.77 show a straight ion tail ∼ 6^{o}.7 long in p.a. 44^{o}, when a red continuum filter (647 nm) showed a fan of dust ∼ 2^{o} long extending in p.a. 10^{o}-40^{o}. Both filters have FWHM = 10 nm. Clouds and knots of material are present along the tail stream on all three nights.

Further total visual magnitude estimates: Nov. 14.75 UT, 5.0 (A. Boattini, Piazzano, Italy, naked eye); 15.70, 5.3 (O. Midtskogen, Tranby, Norway, 0.05-m refractor); 17.78, 4.9 (T. Vanmunster, Landen, Belgium, 0.12-m reflector); 18.69, 5.0 (A. Kereszturi, Budapest, Hungary, 20×60 binoculars); 20.95, 4.9 (D. W. E. Green, Cambridge, MA, 12×50 binoculars); 21.69, 5.5 (H. Dahle, Oslo, Norway, 7×50 binoculars); 22.38, 4.6 (M. Ohkuma, Mt. Fuji, Japan, 10×70 binoculars); 23.74, 5.4 (R. Behrend, La Chaux-de-Fonds, Switzerland, 10×40 binoculars); 25.11, 5.0 (C. S. Morris, La Canada, CA, 10×50 binoculars).

1992 November 27 Daniel W. E. Green